# Permutation invariant matrix quantum thermodynamics

### v2

@@ 1 Introduction
  • We find indications of the features of in-equivalence of ensembles and negative specific heat capacity in the zero-charge sector of a complex one-matrix model (equivalently 2-Hermitian matrix model) having $U(N)$ gauge symmetry. This model has a finite temperature Hagedorn phase transition as $N$ goes to infinity. The negative specific heat capacity, which is the heat capacity rescaled by $1/N^{2}$, at generic fixed temperatures below the Hagedorn temperature vanishes in the large $N$ limit, but remains finite when the temperature is infinitesimally close to the Hagedorn temperature in the large $N$ limit. The evidence again consists of the high temperature limit which is obtained from the path integral combined with computation with group-theoretic formulae for the counting in terms of Littlewood-Richardson coefficients (see eqn (9.29)). The SAGE code is again provided.
@@ 3.1 Super-exponential degeneracies and vanishing Hagedorn temperature
The large $N$ sequence for $S_{N}$ invariants is given by directed multi-graphs with $k$ edges (any number of nodes) [35]. The stable sequence ${\cal Z}(k,\infty)$ grows approximately as $k!$ at large $k$, by using the graph theoretic interpretation. The counting of directed multi-graphs is certainly as large as the counting of undirected graphs. The set of all undirected graphs with fixed number of edges includes the counting with fixed valency types. It is known [45] that the leading large $k$ counting of graphs with $k$ quartic vertices grows as $k\log k$, a fact familiar to quantum field theorists from the factorial growth of Feynman graph counting. Given such a growth of ${\cal Z}(k,\infty)$ as a function of $k$ it is evident that
@@ 3.2.2 Specific heat capacity versus temperature
The data for $x_{{\rm{max}}},a_{{\rm{max}}}$ for a number of sample values of $N$ are as follows :
@@
The data for $x_{{\rm{min}}},a_{{\rm{min}}}$ are as follows :
@@
The value of the specific heat capacity at the maximum, although this is the heat capacity divided by $N^{2}$ (see equation (3.9)), has a peak which grows as $N$ tends to infinity. We find that this maximum value which we denote as $C_{{\rm{sh}};{\rm{max}}}$ is of order $1$ in units of $N\log N$ in the range of $N$ up to $40$, and conjecture that this will be the case as $N\rightarrow\infty$.
@@
Sample data illustrating the evidence which suggests this conjecture is
@@
On the other hand, the value of the specific heat capacity at the minimum, i.e. at $x=x_{{\rm{min}}}$, denoted $C_{{\rm{sh}};{\rm{min}}}$ scales like $x_{c}={\log N\over N}$. The suggests the conjecture.
@@
A sample of the data suggesting this conjecture is
@@
It is also useful to tabulate the energy $U$, evaluated at $x_{min}$ and $x_{max}$. In contrast to $C_{{\rm{sh}};{\rm{max}}}$, and like $C_{{\rm{sh}},{\rm{min}}}$, the natural scale for $U$ in the transition region is, based on evidence from the data, ${\log N\over N}=x_{c}$ In the following table we give the value $U_{{\rm{min}}}\equiv U(x_{{\rm{min}}})$:
@@
Likewise the energy at the maximum $U_{{\rm{max}}}\equiv U(x_{{\rm{max}}})$ is given as:
@@
Figure 4: Energy curve showing location of $x_{{\rm{max}}}$ (in yellow) at the centre of the energy transition region and $x_{{\rm{min}}}$ (in blue) near the high temperature end of the transition. for $N=20$.
@@
Figure 5: Energy per particle and specific heat capacity per particle divided by $N\log N$, plotted together versus temperature $T$ at $N=20$ : illustrating macroscopic scale of the specific heat capacity in the transition region
@@ 3.4 Zeroes in the complex plane
We get an interesting insight into the nature of the transition at $x\sim x_{c}$ by looking at the behaviour of ${\cal Z}(N,x)$ in the complex $x$-plane. See Figures 8 and 9. In both figures, there are a number of zeroes located close to a vertical line near $x=0$. As $N$ increases, the zeroes become closer to each other and to the origin $x=0$. The table below gives the values of the zeroes nearest to the origin, for a range of $N$, illustrating how they approach the origin as $N$ increases. They are tabulated alongside $x_{{\rm{max}}}$ the location on the real $x$-axis of the maximum of the specific heat capacity $C_{{\rm{sh}}}$ in order to illustrate the proximity of these complex zeroes to the real temperature transition point. This proximity becomes sharper as $N$ increases.
@@
Note these zeroes are approaching the origin as $N\rightarrow\infty$. All the poles of ${\cal Z}(N,x)$ are on the unit circle. This is evident from plots and also from the expression 2.13.
@@ 4.2.1 Energy versus micro-canonical temperature
Figure 11 shows a plot of $T_{\rm{micro};\rm{sym}}(k)$ versus the energy $k$. The point $(k=12,T=0.491)$ is at the minimum temperature. For lower temperatures the energy decreases as the energy increaes. For $k\geq 12$, the energy increases with temperature. We will refer to this value of energy where the $T_{\rm{micro}}$ takes a minimum value as $k_{\rm{crit}}$. Using the asymmetric discretisation of the derivative gives the same shape of curve. There are a couple of points at very small $k$ where the variation of $T_{\rm{micro}}$ with $k$ is different from that of $T_{\rm{micro};sym}$ with $k$.
Figure 10: Plot of micro-canonical temperature $k$ at $N=10$ for $k_{max}=40$ using symmetrised derivative – produces consistent negative SHC trend below critical E
@@ 4.3 Equivalence and in-equivalence of ensembles
In Figure 14 we have superposed for $N=15$, a plot of the expectation value of the energy per particle $U$ in the canonical ensemble versus canonical temperature $T$ with a plot of the rescaled per-particle energy level $E={k\over N^{2}}$ versus the micro-canonical temperature. This shows that, as expected, the two plots agree in the high temperature limit but disagree in the low temperature limit. Figure 15 gives the same superposition of plots at $N=20$.
Figure 14: Plot of the expectation value of the energy $U={\cal U}/N^{2}$ in the canonical ensemble versus canonical temperature $T$, superposed upon $E={k\over N^{2}}$ versus identification of $T_{\rm{micro}}$ in micro-canonical ensembles : Equivalence of ensemble above the transition region. This plot is for $N=15$. The micro-canonical data starts at $k=4$ and ends at $k=100$.
@@ 5.1 Cycle structures and the high temperature expansion
For $k=1$,we have ${\rm{Deg}}(N,[1^{N-2},2])=N^{2}-2N+2$. Note that the coefficient of $N$ is $-2B([1^{N-2k},2^{k}])$ where $B$ is the branching number of the partition $[[1^{N-2k},2^{k}]$. The branching number of a general partition $p=[\{a_{i}^{p_{i}}\}]$ is $\sum_{i}p_{i}(a_{i}-1)$. It has an interpretation in terms of branched covers and plays a role in the Riemann-Hurwitz formula for the genus of covering surfaces in two dimensions. We will find shortly that this property, that the coefficient of $N$ in the degree is the branching number of $p$, holds for general $p$.
@@
Developing further the applications of $x_{c}$ and $x_{\rm{bkdn}}$, we present Figure 17 where we use the data from Tables 1 and 3.2.2 and fit the corresponding critical value of $x$ against $1/N$. The data shows that the critical temperatures are driven to zero as expected and established by the high temperature analysis. The figure shows that the microcanonical transition is consistently at a lower temperature than the canonical one as expected from Figure 15 but both go to zero at infinite $N$ as expected from the known stable structure of the degeneracies and from our high temperature analysis which establishes that the transition has to occur for $x\lessapprox x_{c}=\frac{\log N}{N}$. Our fitting curves are motivated by the functional form suggested by the first three terms in LABEL:xbkform.
@@ 8 Hagedorn transitions in $U(N)$ invariant models and negative SHC in tensor models
Figure 18: Micro-canonical energy versus temperature for 3-index tensor at $N=3$ with $k$ equals $3$ to $13$ using the symmetric $D_{\rm{sym}}$ discrete derivative
Figure 19: Micro-canonical energy versus temperature for $3$-index tensor $N=4$ with $k$ equals $3$ to $12$ using the symmetric $D_{\rm{sym}}$ discrete derivative. Note the curve turns around, i.e. SHC become positive at higher energy.
The data ${\cal Z}(n,N)$ is copied to Mathematica and used to plot the micro-canonical energy $n$ versus the micro-canonical tempertature. The cases $N=3,4$ are shown in Figures 18 and 19. They demonstrate the turn-over from negative SHC to positive SHC. In the GPIMQM case, with the more explicit formula (2.13) in hand we were able to demonstrate this behaviour for $N$ up to $40$. Producing the data at higher $N$ and providing detailed evidence that this system is, as argued here, in the same universality class as the GPIMQM is an interesting challenge.
@@ 9 Negative SHCs in matrix quantum mechanical models and AdS/CFT
The micro-canonical inverse temperature
$\beta_{\rm{micro}}=T_{\rm{micro}}^{-1}={\partial S(k)\over\partial k}=abk^{b-1}$ (9.2)
$T_{\rm{micro}}=a^{-1}b^{-1}k^{1-b}$ (9.3)
${\partial T_{\rm{micro}}\over\partial k}=a^{-1}b^{-1}(1-b)k^{-b}$ (9.4)
$C_{{\rm{sh}};\rm{micro}}=\left({\partial T_{\rm{micro}}\over\partial k}\right)^{-1}={ab\over(1-b)}k^{b}$ (9.5)
@@
We consider
@@
$\beta_{\rm{micro}}(k)=ab\log^{b-1}(k)+a\log^{b}(k)$ (9.7)
${\partial\beta_{\rm{micro}}\over\partial k}=abk^{-1}\log^{b-2}(k)(b+\log(k)-1)$ (9.8)
@@
the micro-SHC is
$\displaystyle C_{\rm{hc};\rm{micro}}(k)=-\frac{ak\log^{b}(\log(k))(b+\log(k)\log(\log(k)))^{2}}{b(b+(\log(k)-1)\log(\log(k))-1)}$ (9.12)
and at large $k$,
$\displaystyle C_{\rm{hc};\rm{micro}}(k)\sim{-ak\over b}(\log(\log(k))^{b+1}\log(k)$ (9.13)
At large integer $k$, we get negative SHC for the super-exponential growths $a>0,b>0$ while for the near-exponential but sub-exponential growths, we have positive SHC.
@@
In the large $k$ limit, which is relevant to convergence of the partition function
$\displaystyle\beta_{\rm{micro}}(k\rightarrow\infty)=\log d$ (9.16)
The derivative
$\displaystyle{\partial\beta_{\rm{micro}}(k)\over\partial k}=-{a\over k^{2}}$ (9.17)
The micro-canonical heat capacity ${}_{C}{\rm{hc};\rm{micro}}(k)$ is
$\displaystyle C_{\rm{hc};\rm{micro}}(k)$ $=$ ${k^{2}\over a}(\log d+{a\over k})^{2}$ (9.18)
In the large $k$ limit,
@@ 9.2 Negative SHCs from zero-charge sector of complex matrix quantum mechanics and AdS/CFT
which specialises (9.28) to $m=n$. The output for $n=1$ to $n=13$ is
@@ 10 Summary and Outlook
In our discussion of negative SHC in the zero charge sector of a complex matrix (or equivalently the Hermitian 2-matrix model) we discussed the interpretation in terms of branes and anti-branes. A derivation of this feature directly from brane/anti-brane systems in the bulk of AdS would be fascinating. Existing general discussions, e.g. [70][71], of negative specific heat heat capacities in connection with long-range interactions are likely to be useful as ingredients in such a project.

### v3

@@ 1 Introduction
  • We find indications of the features of in-equivalence of ensembles and negative specific heat capacity in the zero-charge sector of a complex one-matrix model (equivalently 2-Hermitian matrix model) having $U(N)\times U(1)$ gauge symmetry. This model has a finite temperature Hagedorn phase transition as $N$ goes to infinity. The negative specific heat capacity, which is the heat capacity rescaled by $1/N^{2}$, at generic fixed temperatures below the Hagedorn temperature vanishes in the large $N$ limit, but remains finite when the temperature is infinitesimally close to the Hagedorn temperature in the large $N$ limit. The evidence again consists of the high temperature limit which is obtained from the path integral combined with computation with group-theoretic formulae for the counting in terms of Littlewood-Richardson coefficients (see eqn (9.20)). The SAGE code is again provided.
@@ 3.1 Super-exponential degeneracies and vanishing Hagedorn temperature
The large $N$ sequence for $S_{N}$ invariants is given by directed multi-graphs with $k$ edges (any number of nodes) [35]. The stable sequence ${\cal Z}(k,\infty)$ grows approximately as $k!$ at large $k$, by using the graph theoretic interpretation. The counting of directed multi-graphs is certainly as large as the counting of undirected graphs. The set of all undirected graphs with fixed number of edges includes the counting with fixed valency types. It is known [45] that the leading large $k$ counting of vacuum graphs with $k$ quartic vertices grows as ${(4k)!\over(2k)!(k!)2^{2k}(4!)^{k}}$, the logarithm of which grows as $k\log k$ in the large $k$ limit. This factorial growth of Feynman graph counting is well-known to quantum field theorists. Given such a growth of ${\cal Z}(k,\infty)$ as a function of $k$ it is evident that
@@ 3.2.2 Specific heat capacity versus temperature
The data for $x_{{\rm{max}}},a_{{\rm{max}}}$ for a number of sample values of $N$ are given in Table 1.
Table 1: Location of maximum of SHC
@@
The data for $x_{{\rm{min}}},a_{{\rm{min}}}$ are given in Table 2.
Table 2: Location of minimum of SHC
@@
At high temperatures, the specific heat capacity matches that of $N^{2}$ simple harmonic oscillators, as we derive in section 5. Using (A.3), it follows that the specific heat capacity has positive gradient as it approaches the limiting value at $x=1$. For $x>x_{{\rm{max}}}$, there is therefore at least one minimum. The data, with the range of values of $N$ considered, suggests that it is a general property valid for generic $N$ that there is exactly one minimum. Since $a_{{\rm{min}}}$ in Table 2 is slowly increasing, we expect that $x_{{\rm{min}}}$ approaches 00 at least as fast as ${\log N\over N}$ as $N\rightarrow\infty$. The location $x_{{\rm{max}}}$ of the maximum, which is less than $x_{{\rm{min}}}$, therefore also approaches zero at least as rapidly as ${\log N\over N}$. This is in line with the estimate of ${\log N\over N}$ for the transition region based on the breakdown of the high-temperature expansion (section 5.2).
The value of the specific heat capacity at the maximum, although this is the heat capacity divided by $N^{2}$ (see equation (3.9)), has a peak which grows as $N$ tends to infinity. We find that this maximum value which we denote as $C_{{\rm{sh}};{\rm{max}}}$ is of order $1$ in units of $N\log N$ in the range of $N$ up to $40$, and conjecture (Conjecture 1) that this will be the case as $N\rightarrow\infty$.
@@
Sample data illustrating the evidence which suggests this conjecture is in Table 3.
Table 3: ${C_{{\rm{sh}};{\rm{max}}}\over N\log N}$ as a function of $N$
@@
On the other hand, the value of the specific heat capacity at the minimum, i.e. at $x=x_{{\rm{min}}}$, denoted $C_{{\rm{sh}};{\rm{min}}}$ scales like $x_{c}={\log N\over N}$. The suggests the Conjecture 2 which follows.
@@
A sample of the data suggesting this conjecture is in Table 4.
Table 4: ${C_{{\rm{sh}};{\rm{min}}}\over x_{c}}$ as a function of $N$
@@
It is also useful to tabulate the energy $U$, evaluated at $x_{min}$ and $x_{max}$. In contrast to $C_{{\rm{sh}};{\rm{max}}}$, and like $C_{{\rm{sh}},{\rm{min}}}$, the natural scale for $U$ in the transition region is, based on evidence from the data, ${\log N\over N}=x_{c}$. In the Table 5 we give the value $U_{{\rm{min}}}\equiv U(x_{{\rm{min}}})$:
Table 5: ${U_{{\rm{min}}}\over x_{c}}$ as a function of $N$
@@
Likewise the energy at the maximum $U_{{\rm{max}}}\equiv U(x_{{\rm{max}}})$ is given in Table 6.
Table 6: ${U_{{\rm{max}}}\over x_{c}}$ as a function of $N$
@@
Figure 4: Energy curve showing location of $x_{{\rm{max}}}$ (in yellow) at the centre of the energy transition region and $x_{{\rm{min}}}$ (in green) near the high temperature end of the transition. for $N=20$.
@@ 3.4 Zeroes in the complex plane
The coalescence of zeroes in the complex plane in thermodynamic limits has long been recognised as a signature of phase transitions in the statistical physics literature [48, 49, 50].
In the case of the finite temperature large $N$ Hagedorn transition of the 2-matrix model, the analysis of the singularity structure of the partition function at finite $N$, in the complex $x=e^{-1/T}$ plane, has been found to be illuminating [10]. The location of $x_{c}=e^{-1/T_{c}}$ of the $N\rightarrow\infty$ phase transition has been found to be the convergence point of two arcs of zeroes in the complex plane (Fig. 1 of [10]). In the present case of a Hagedorn temperature which tends to zero as $N\rightarrow\infty$, we find a similar behaviour of zeroes in the complex $x$-plane, in the neighbourhood of the location of the peak in the specific heat capacity along the real $x$-axis. This is illustrated in Figures 7 and 8. In both figures, there are a number of zeroes located close to a vertical line near $x=0$. As $N$ increases, the zeroes become closer to each other and to the origin $x=0$. Table 7 below gives the values of the zeroes nearest to the origin, for a range of $N$, illustrating how they approach the origin as $N$ increases. They are tabulated alongside $x_{{\rm{max}}}$ the location on the real $x$-axis of the maximum of the specific heat capacity $C_{{\rm{sh}}}$ in order to illustrate the proximity of these complex zeroes to the real temperature transition point. This proximity becomes sharper as $N$ increases.
Table 7: complex plane zeroes and $x_{{\rm{max}}}$
@@
Note that these zeroes are approaching the origin as $N\rightarrow\infty$. All the poles of ${\cal Z}(N,x)$ are on the unit circle. This is evident from plots and also from the expression (2.13).
@@ 4.2.1 Energy versus micro-canonical temperature
Figure 10 shows a plot of $T_{\rm{micro};\rm{sym}}(k)$ versus the energy $k$. The point $(k=12,T=0.491)$ is at the minimum temperature. In the lower energy branch of the curve, the energy decreases as the temperature increases. For $k\geq 12$, the energy increases with temperature. We will refer to this value of energy where the $T_{\rm{micro}}$ takes a minimum value as $k_{\rm{crit}}$. Using the asymmetric discretisation of the derivative gives the same shape of curve. There are a couple of points at very small $k$ where the variation of $T_{\rm{micro}}$ with $k$ is different from that of $T_{\rm{micro};sym}$ with $k$.
Figure 9: Plot of micro-canonical temperature versus $k$ at $N=10$ for $k_{max}=40$ using symmetrised derivative – produces consistent negative SHC trend below critical E
@@ 4.3 Equivalence and in-equivalence of ensembles
In Figure 13 we have superposed for $N=20$, a plot of the expectation value of the energy per particle $U$ in the canonical ensemble versus canonical temperature $T$ with a plot of the rescaled per-particle energy level $E={k\over N^{2}}$ versus the micro-canonical temperature. This shows that, as expected, the two plots agree in the high temperature limit but disagree in the low temperature limit. The same features hold for different values of $N$.
@@ 5.1 Cycle structures and the high temperature expansion
For $k=1$,we have ${\rm{Deg}}(N,[1^{N-2},2])=N^{2}-2N+2$. Note that the coefficient of $N$ is $-2B([1^{N-2k},2^{k}])$ where $B$ is the branching number of the partition $[1^{N-2k},2^{k}]$. The branching number of a general partition $p=[\{a_{i}^{p_{i}}\}]$ is $\sum_{i}p_{i}(a_{i}-1)$. It has an interpretation in terms of branched covers and plays a role in the Riemann-Hurwitz formula for the genus of covering surfaces in two dimensions. We will find shortly that this property, that the coefficient of $N$ in the degree is the branching number of $p$, holds for general $p$.
@@
Developing further the applications of $x_{c}$ and $x_{\rm{bkdn}}$, we present Figure 15 where we use the data from Tables 8 and 3 and fit the corresponding critical value of $x$ against $1/N$. The data shows that the critical temperatures are driven to zero as expected and established by the high temperature analysis. The figure shows that the microcanonical transition is consistently at a lower temperature than the canonical one as expected from Figure 13 but both go to zero at infinite $N$ as expected from the known stable structure of the degeneracies and from our high temperature analysis which establishes that the transition has to occur for $x\lessapprox x_{c}=\frac{\log N}{N}$. Our fitting curves are motivated by the functional form suggested by the first three terms in LABEL:xbkform.
@@ 8 Hagedorn transitions in $U(N)$ invariant models and negative SHC in tensor models
Figure 16: Micro-canonical energy versus temperature for 3-index tensor at $N=3$ with $k$ equals $4$ to $14$ using the symmetric $D_{\rm{sym}}$ discrete derivative
Figure 17: Micro-canonical energy versus temperature for $3$-index tensor $N=4$ with $k$ equals $4$ to $13$ using the symmetric $D_{\rm{sym}}$ discrete derivative. Note the curve turns around, i.e. SHC become positive at higher energy.
The data ${\cal Z}(n,N)$ is copied to Mathematica and used to plot the micro-canonical energy $n$ versus the micro-canonical tempertature. The cases $N=3,4$ are shown in Figures 16 and 17. They demonstrate the turn-over from negative SHC to positive SHC. In the GPIMQM case, with the more explicit formula (2.13) in hand we were able to demonstrate this behaviour for $N$ up to $40$. Producing the data at higher $N$ and providing detailed evidence that this system is, as argued here, in the same universality class as the GPIMQM is an interesting challenge.
@@ 9 Negative SHCs in matrix quantum mechanical models and AdS/CFT
The micro-canonical inverse temperature is
$\beta_{\rm{micro}}=T_{\rm{micro}}^{-1}={\partial S(k)\over\partial k}$ (9.2)
which leads to
$C_{\rm{hc};\rm{micro}}=\left({\partial T_{\rm{micro}}\over\partial k}\right)^{-1}={ab\over(1-b)}k^{b}$ (9.3)
@@
We consider weakly super-exponential degeneracies of the form
@@
we obtain negative SHC for the super-exponential growths $a>0,b>0$ while for the near-exponential but sub-exponential growths, we have positive SHC.
@@
In the large $k$ limit, which is relevant to the convergence of the partition function, the micro-canonical heat capacity is calculated to be:
@@
To summarise, examples of where the degeneracy $\Omega(k)$ as a function of energy $k$ in the micro-canonical ensemble fails to be concave, leading to negative heat capacity are given below in Table 9
Table 9: Degeneracies and negative heat capacities
$\Omega(k)$ Parameter ranges Description
$e^{ak^{b}}$ $a>0,b>1$ Super-exponential
$e^{ak(\log k)^{b}}$ $a>0,b>0$ Weakly super-exponential
$e^{ak(\log(\log k))^{b}}$ $a>0,b>0$ Weakly super-exponential
$e^{k\log d}k^{a}$ $d>1,a<0$ Sub-exponential (power-law corrected exponential)
@@ 9.2 Negative SHCs from zero-charge sector of complex matrix quantum mechanics and AdS/CFT
which specialises (9.19) to $m=n$. The output for $n=1$ to $n=18$ is
@@ 10 Summary and Outlook
In our discussion of negative SHC in the zero charge sector of a complex matrix (or equivalently the Hermitian 2-matrix model) we discussed the interpretation in terms of branes and anti-branes. A derivation of this feature directly from brane/anti-brane systems in the bulk of AdS would be fascinating. Existing general discussions, e.g. [73][74], of negative specific heat capacities in connection with long-range interactions are likely to be useful as ingredients in such a project.